If the solutions to a quadratic equation are 2 and - 4/3, what is the equation?
Write the quadratic in standard form: Ax2 + Bx + C = 0.
A, B, and C must be integers - no fractions, no decimals - and A must be positive

Answer by Alan3354(69443)

If the solutions to a quadratic equation are 2 and - 4/3, what is the equation?
Write the quadratic in standard form: Ax2 + Bx + C = 0.
===========
(x-2)*(x+4/3) = 0
(x-2)*(3x+4) = 0
3x^2 - 2x - 8 = 0 is AN equation with the smallest integer coefficients.
---------
n*(3x^2 - 2x - 8) = 0 where n is any real number is also an equation.